OpenStudy (anonymous):

Use the substitution method to find the solution(s) to the system of equations. Show all work for full credit. y = –x^2 y = x

OpenStudy (anonymous):

\[ y=-x^2\\ y=x\\ \to\\ x=-x^2 \]

OpenStudy (anonymous):

\[ x^2+x=0\\ x(x+1)=0\\ x=0,x=-1 \]

OpenStudy (anonymous):

Thus the two solutions are\[ (0,0),(-1,-1) \]

OpenStudy (anonymous):

Prodigious!

OpenStudy (anonymous):